# LOCKET: Robust Feature-Locking Technique for Language Models

### v1

@@ 5 Experimental Setup
Datasets. We use four datasets, each corresponding to a premium feature: (i) Math (M)which contains challenging math problems Hendrycks et al. (2021b); (ii) SQL (Q)for structured query language generation Zhong et al. (2017); Yu et al. (2018); (iii) Text Summarization (S)from the SAMSum dataset Gliwa et al. (2019); (iv) General Knowledge (U)from the MMLU benchmark Hendrycks et al. (2021a). For $D_{\text{auth}}$, we use samples from the UltraChat dataset Ding et al. (2023). We describe details about the datasets in Appendix B.
Models. Following Greenblatt et al. (2024a), we use two LLMs: DeepSeek-7B-Math (specialized in math) Shao et al. (2024), and DeepSeek-7B-Coder (specialized in coding) Guo et al. (2024). We additionally use one general-purpose conversation model, Llama-3-8B-Instruct Grattafiori et al. (2024). DeepSeek-7B-Coder, trained on code, includes a system prompt that refuses non-coding queries. Unlike other models, we evaluate DeepSeek-7B-Coder on coding tasks. We describe fine-tuning hyperparameters in Appendix B. Evaluations are conducted with temperature set to zero.
@@
Comparison with Prior Work. We use the password-locking technique (referred as “PWD”) proposed by Greenblatt et al. (2024b), where prompts with correct password produce a useful response. Instead of locking by generating a useless response Greenblatt et al. (2024b), we fine-tune for refusal which resembles Tang et al. (2024).
@@ 6.1 Evaluation of LOCKET Merging
We compare LOCKET merging with existing merging methods in terms of their impact on utility-preservation. We unlock one feature and lock all remaining features (to capture the worst-case impact) by merging the corresponding adapters into the base LLM. We then measure the unlocked feature’s refusal rate ($100-\text{utility}$) to determine if the merging causes over-refusal on the unlocked feature. Figure 2 (Top) shows that LOCKET yields significantly lower refusal rates (i.e., higher utility) than other approaches.
Figure 2: These are illustrative examples for DeepSeek-7B-Math. We observe similar patterns in other models and locking combinations. (Top) LOCKET merging significantly reduces U’s over-refusal rate compared to prior merging methods; (Bottom) Scaling hyperparameter $\tau$ should be chosen to balance trade-off between effectiveness and utility. Here, only U is unlocked; vertical line indicates the sweet spot for $\tau$ (high refusal for locked and no utility drop on unlocked features). See Appendix B for other $\tau$ values.
Selecting the Scaling Hyperparameter $\tau$. To show the trade-offs from varying $\tau$, we illustrate using DeepSeek-7B-Math by locking three features (M, Q, and S), and leaving U as unlocked (Figure 2: Bottom). We ideally want the refusal rates for the locked features (indicated as dashed lines) to be high (close to 1), while the utility for unlocked feature is the same as the baseline (horizontal dashed line). We find that the value of $\tau=0.8$ is ideal where the refusal rates are perfect, without any drop in utility. We use the same hyperparameter tuning approach for selecting $\tau$, to lock other features and their combinations (Appendix B).
@@ 6.2 R1 (Effectiveness)
Results for LOCKET. We evaluate effectiveness of feature locking by measuring the utility wrt. the locked feature. Ideally, this is zero (indicating 100% refusal rate). Table 2 shows the results, where effective locking is in blue, and ineffective locking in orange. For perfect effectiveness, the diagonal corresponding to the same feature in the row and column, should be zero (or blue). This is indeed the case, suggesting LOCKET’s effectiveness.
Table 2: LOCKET is effective and utility-preserving: “Baseline” is the original model behaviour without FLoTE. For effectiveness (R1), we use blue to indicate complete locking and orange otherwise. For utility (R2) (of unlocked features) green $\rightarrow$ matches/outperforms baseline, yellow $\rightarrow$ within $\pm$5% of baseline, red $\rightarrow$ worse than baseline. Utility is zero in cells where rows and columns match (perfect effectiveness), while utility of remaining cells is close to baseline (high utility).
@@
Comparison with PWD. For illustrating the comparison of LOCKET and PWD, we use DeepSeek-7B-Math with “Math” (M) locked. For perfect effectiveness, we want the utility of the same feature (i.e., M) to be zero. This is indeed the case for both PWD and LOCKET, indicating perfect effectiveness.
@@ 6.3 R2 (Utility-Preserving)
Results for LOCKET. Table 2 also shows the utility of LOCKET. We evaluate utility preservation by measuring performance on unlocked features relative to the baseline. For perfect utility, the value in non-diagonal cells (where features in rows and columns do not match), should match baseline values. We use green to indicate outperforming/matching baseline, yellow within $\pm$5% of baseline, and red worse than baseline. When locking single features, LOCKET can successfully preserve the utility of unlocked features in two of the three rows (green). Locking M (both DeepSeek-7B-Math and Llama-3-8B-Instruct) and Q (for DeepSeek-7B-Math) results in a small utility drop of 2-3% in U (yellow), while locking U in Llama-3-8B-Instruct results in a 6% drop in M (red). This is from interference among features due to the presence of math-related questions in U (discussed in §7).
Comparison with PWD. For illustrating the comparison of LOCKET and PWD, we use DeepSeek-7B-Math with “Math” (M) locked. Ideally, we want the utility of the unlocked feature (i.e., Q, S, and U) to be similar to the baseline (original model behavior without locking) in Table 2. Evaluating Q, S, and U, LOCKET matches baseline utility for Q and S, with a minor 2% drop on U (due to M interference). PWD shows a significant 12% drop on S, and a similar 2% drop on U. This indicates LOCKET better preserves utility, likely because it augments specific layers of the frozen LLM, whereas PWD fine-tunes and overwrites the original weights.
6.4 R3 (Robustness)
Table 3: LOCKET is more robust than PWD: Attack success rates (ASR) on adversarial prompts where lower score is better. Results are written as “<PWD> | <LOCKET >” with green indicating LOCKET outperforms PWD, red for worse, and yellow for similar (within $\pm$5%).
@@ 6.4 R3 (Robustness)
Adversarial prompts are transferable across models Zou et al. (2023); Mehrotra et al. (2024). Specifically, prompts designed to evade a model with a single feature locked (e.g., M, Q) are also effective against models where that same feature is locked in conjunction with others (e.g., M + S, Q + U). Therefore, to measure the upper bound on robustness, we use AdptJB to evaluate models with a single feature locked, and consider the following attacks: Many-shot Anil et al. (2024), GCG Zou et al. (2023), and AutoDAN-Turbo Liu et al. (2025) (hyperparameters in Appendix B). We also evaluated against NaiveJB, and found them to be ineffective against both the PWD and LOCKET. Hence we omit those results. Table 3 shows the ASRs in the format “PWD | LOCKET”. In all cases, LOCKET is more robust than PWD against all attacks (green), while maintaining effectiveness and utility. Unlike prior work Greenblatt et al. (2024a); Tang et al. (2024), since LOCKET does not use any secret credentials, it is protected against credential stealing and unauthorized redistribution (R3.4).
@@ 6.5 R4 (Scalable)
To demonstrate scalability of LOCKET and compare with PWD, we evaluate effectiveness (R1) and utility (R2) on locking multiple features. Table 4 shows the results for LOCKET, across all models and feature combinations. The color coding is the same as in Table 2. In all cases, LOCKET is perfectly effective (blue) with some utility drop ($\leq$7%), primarily due to interference between M and U. Table 5 compares LOCKET and PWD with two or three features locked (M + Q, M + Q + S) for DeepSeek-7B-Math, we observe similar behavior for other models as well (Appendix C). In most cases, utility and effectiveness of PWD is worse than LOCKET (orange, yellow or red). This suggests that LOCKET scales to more than two features, unlike PWD. PWD’s full fine-tuning likely causes “catastrophic forgetting” Kotha et al. (2024), where training refusal for one feature harms others.
Table 5: Comparison of LOCKET with prior work: Scalability w.r.t. R1 and R2 of LOCKET with prior work (“PWD”) Greenblatt et al. (2024a); Tang et al. (2024) locking DeepSeek-7B-Math (more in Appendix C). Color coding for scalability, are same as Table 2.
@@ 7 Discussions and Summary
Other Applications. While we focus on using FLoTEs for pay-to-unlock schemes, they have other potential applications. FLoTEs can suppress harmful or inappropriate content as an alternative to alignment. They also offer a robust feature-level unlearning method by locking behaviors. Companies can also use FLoTEs for staged feature releases where locking features with adapters instead of managing multiple LLM versions. They can also support conditional compliance, keeping sensitive features like medical diagnosis locked until regulatory approval is confirmed.
Arms Race for Robustness. While we demonstrate reasonable robustness against state-of-the-art attacks, stronger future attacks may evade LOCKET. Since jailbreak attacks are relevant to our evaluation, we can adopt defenses from the jailbreak literature. For such attacks, LOCKET ’s adapters can be fine-tuned to maintain robustness.
Feature interference. Ideally, the features should be non-overlapping, but in practice, there could be interference between features. We only observed interference in a few cases, but as part of future work, we intend to explore the design of FLoTEs which are resistant to interference.
Summary. We identify a new application of pay-to-unlock features in LLMs. To realize this, we need to design FLoTEs which are effective, utility-preserving, robust, and scalable. None of the prior work meets all requirements. We propose LOCKET, the first robust and scalable FLoTE.
@@
We identify a couple of limitations, for future work:
  • Model Types. We only consider three model types in our work following prior work on password-locking Greenblatt et al. (2024a). We leave a more comprehensive evaluation across other model types as future work.
  • Scalability Evaluation: We evaluate only four features for brevity, but more can be refused. Prior work Lee et al. (2025) shows that around 12 adapters can be merged with at most $20\%$ utility drop. Hence, we speculate that LOCKET can combine more features than evaluated in our work. An empirical validation is left for future.
  • Impact of Computation/Energy Cost. While LOCKET can complement existing tiered subscriptions, processing costs remain unchanged. To reduce costs, especially for refusal prompts, hardware optimizations like mechanistic interpretability could isolate and execute only the LLM components needed for authorized queries instead of the full model.
@@ Appendix A Notations
Appendix B Implementation Details
@@
Computational Cost. Experiments are done using 8 * NVIDIA A100 GPUs which consume in total round 5000 GPU hours.
Appendix C Scalability Comparison with PWD

### v2

@@ 5 Experimental Setup
Datasets. We use four datasets, each corresponding to a premium feature: (i) Math (M)which contains challenging math problems Hendrycks et al. (2021b); (ii) SQL (Q)for structured query language generation Zhong et al. (2017); Yu et al. (2018); (iii) Text Summarization (S)from the SAMSum dataset Gliwa et al. (2019); (iv) General Knowledge (U)from the MMLU benchmark Hendrycks et al. (2021a). For $D_{\text{auth}}$, we use samples from the UltraChat dataset Ding et al. (2023). We describe details about the datasets in Appendix E.
Models. Following Greenblatt et al. (2024a), we use two LLMs: DeepSeek-7B-Math (specialized in math) Shao et al. (2024), and DeepSeek-7B-Coder (specialized in coding) Guo et al. (2024). We additionally use one general-purpose conversation model, Llama-3-8B-Instruct Grattafiori et al. (2024). DeepSeek-7B-Coder, trained on code, includes a system prompt that refuses non-coding queries. Unlike other models, we evaluate DeepSeek-7B-Coder on coding tasks. We describe fine-tuning hyperparameters in Appendix E. Evaluations are conducted with temperature set to zero.
@@
Comparison with Prior Work. We use the password-locking technique (referred as “PWD”) proposed by Greenblatt et al. (2024b) and used in related works (§2), where prompts with correct password produce a useful response. Instead of locking by generating a useless response Greenblatt et al. (2024b), we fine-tune for refusal which resembles Tang et al. (2024).
@@ 6 Evaluation
Figure 2: Illustrative examples for DeepSeek-7B-Math: We observe similar patterns in other models and locking combinations. (Top) LOCKET merging significantly reduces U’s over-refusal rate compared to prior merging methods. (Bottom) The scaling hyperparameter $\tau$ should be chosen to balance the trade-off between effectiveness and utility. Here, only U is unlocked; the vertical line indicates the sweet spot for $\tau$ (high refusal for locked and no utility drop on unlocked features). See Appendix E for other $\tau$ values.
@@ 6.1 Evaluation of LOCKET Merging
We compare LOCKET merging with existing merging methods in terms of their impact on utility-preservation. We unlock one feature and lock all remaining (to capture the worst-case impact) by merging and attaching the corresponding adapters to the base LLM. We then measure the unlocked feature’s refusal rate ($100-\text{utility}$) to determine if the merging causes over-refusal on the unlocked feature. Figure 2 (Top) shows that LOCKET yields significantly lower refusal rates than the others (detailed results in Appendix G)
Selecting the Scaling Hyperparameter $\tau$. To show the trade-offs from varying $\tau$, we illustrate using DeepSeek-7B-Math by locking three features (M, Q, and S), and leaving U as unlocked (Figure 2: Bottom). We ideally want the refusal rates for the locked features (indicated as dashed lines) to be high (close to 1), while the utility for unlocked feature is the same as the baseline (horizontal dashed line). We find that the value of $\tau=0.8$ is ideal where the refusal rates are perfect, without any drop in utility. We use the same hyperparameter tuning approach for selecting $\tau$, to lock other features and their combinations (Appendix E).
@@ 6.2 R1 (Effectiveness)
Results for LOCKET. We evaluate effectiveness by measuring utility on locked features, which should ideally be zero (100% refusal rate). Table 2 shows the results, with effective locking in blue. For perfect effectiveness, diagonal cells (same feature in row and column) should be zero. This holds across all cases, confirming LOCKET’s effectiveness.
Table 2: LOCKET is effective and utility-preserving: “Baseline” is the original model behaviour without FLoTE. For effectiveness (R1), we use blue to indicate complete locking. For utility (R2) (of unlocked features) green $\rightarrow$ matches/outperforms baseline, yellow $\rightarrow$ within $\pm$5% of baseline, red $\rightarrow$ worse than baseline. Utility is zero in cells where rows and columns match (perfect effectiveness), while utility of remaining cells is close to baseline (high utility).
@@
Comparison with PWD. Using DeepSeek-7B-Math with “Math” (M) locked, both PWD and LOCKET achieve zero utility on M, indicating perfect effectiveness.
@@ 6.3 R2 (Utility-Preserving)
Results for LOCKET. Table 2 also shows utility preservation, measured by performance on unlocked features relative to baseline. Non-diagonal cells should match baseline for perfect utility. We use green for matching/outperforming baseline, yellow for within $\pm$5%, and red for worse. When locking single features, LOCKET preserves utility in two of three rows (green). Locking M (both DeepSeek-7B-Math and Llama-3-8B-Instruct) or Q (DeepSeek-7B-Math) causes a small 2-3% drop in U (yellow), while locking U in Llama-3-8B-Instruct causes a 6% drop in M (red) due to feature interference from math-related questions in U7).
Comparison with PWD. Using DeepSeek-7B-Math with M locked, LOCKET matches baseline utility for Q and S, with only a 2% drop on U (due to interference). In contrast, PWD shows a significant 12% drop on S and similar 2% on U. LOCKET better preserves utility because it augments specific layers of a frozen LLM, whereas PWD fine-tunes and overwrites original weights.
Table 3: LOCKET is more robust than PWD: Attack success rates (ASR) on adversarial prompts (lower is better). Results are written as “<PWD> | <LOCKET >” with green indicating LOCKET outperforms PWD, red for worse, and yellow for similar (within $\pm$5%).
@@
6.4 R3 (Robustness)
Adversarial prompts transfer across models Zou et al. (2023); Mehrotra et al. (2024): prompts evading a model with one locked feature (e.g., M) also work when that feature is locked with others (e.g., M +S). To measure upper-bound robustness, we attack single-feature-locked models with Many-shot Anil et al. (2024), GCG Zou et al. (2023), and AutoDAN-Turbo Liu et al. (2025) (hyperparameters in Appendix E). NaiveJB omitted as they were ineffective against both PWD and LOCKET.
Table 3 shows ASRs as “PWD | LOCKET”. LOCKET is consistently more robust than PWD (green) while maintaining effectiveness and utility. Since LOCKET uses no secret credentials, it is also protected against credential stealing and redistribution (R3.4Tang et al. (2024).
@@ 6.5 R4 (Scalable)
We evaluate scalability by measuring effectiveness (R1) and utility (R2) when locking multiple features. Table 4 shows results across all models and feature combinations, with color coding as in Table 2. LOCKET achieves perfect effectiveness (blue) with utility drops $\leq$7% from M-U interference.
Table 5 compares LOCKET and PWD with two or three features locked (M + Q, M + Q + S) for DeepSeek-7B-Math; other models show similar patterns (Appendix H). PWD’s utility and effectiveness are mostly worse (orange, yellow, or red), suggesting LOCKET scales better. PWD’s full fine-tuning likely causes “catastrophic forgetting” Kotha et al. (2024), where training refusal for one feature harms others.
Table 5: Comparison of LOCKET with prior work: Scalability w.r.t. R1 and R2 of LOCKET with prior work (“PWD”) Greenblatt et al. (2024a); Tang et al. (2024) locking DeepSeek-7B-Math (more in Appendix H). Color coding same as Table 4. For R1, we use blue to indicate complete locking and orange otherwise.
@@ 7 Discussions and Summary
Overhead of LOCKET. We evaluate the runtime overhead of LOCKET on a single A100 40GB PCIe GPU, averaged over 5 runs. Attaching an adapter takes $1\pm 0.06$ second, while detaching takes $0.02\pm 0.00$ seconds. Attaching only needs to occur once per user session (i.e., at login), and inference latency and throughput remain unaffected: Time-to-First-Token (TTFT) is $3\pm 0.30$ms regardless of the number of adapters attached. These results show that serving costs for low-tier and paying users are comparable, making pay-to-unlock economically feasible. Storage size of individual LOCKET adapters is also a fraction of the base model size. A single locket adapter is only 1.6 - 1.7% of the base model’s total parameter count (120M for DeepSeek-7B-Math/DeepSeek-7B-Coder, 130M for Llama-3-8B-Instruct). Overall, LOCKET incurs minimal overhead in both storage and computation.
Model Types. Following prior work on password-locking Greenblatt et al. (2024a), we consider three model types (§6). To validate that our adapter-based feature-locking technique can scale effectively to larger models, we provide evaluation results on Llama-3-70B-Instruct in Appendix H. We found that LOCKET remains effective and utility-preserving (other than expected interference between M and U). We leave a more comprehensive evaluation across other models as future work.
Scalability Evaluation. We considered four features for the main evaluation, though our approach naturally extends to more. Prior work Lee et al. (2025) demonstrates that roughly 8 adapters can be merged with at most 15% utility drop. Our findings align with this trend: in Appendix I, we show for DeepSeek-7B-Math that LOCKET can scale to 8 features while keeping utility and effectiveness degradation within the 15% bound established in prior work. We leave further scalability improvements as future work.
Summary. We identify pay-to-unlock features as a new LLM application requiring FLoTEs that are effective, utility-preserving, robust, and scalable. No prior work meets all requirements. We propose LOCKET, the first robust and scalable FLoTE.
@@
  • Feature Interference. Ideally, features should be non-overlapping, but interference can occur in practice. We observed this in a few cases (Math and MMLU, which contains math-related questions); designing interference-resistant FLoTEs remains future work. Note that this interference is distinct from catastrophic forgetting, a failure mode of full fine-tuning methods like PWD where locking one feature overwrites knowledge of others. LOCKET does not suffer from forgetting because adapters attach dynamically to a frozen base model without permanently modifying its weights. We intend to explore potential mitigation strategies such as pre-processing feature datasets to remove semantically overlapping samples before adapter training in follow-up work.
  • Arms Race for Robustness. While we demonstrate robustness against state-of-the-art attacks, stronger future attacks may evade LOCKET. Since jailbreak attacks are relevant, defenses from that literature can be adopted, and LOCKET’s adapters can be fine-tuned accordingly.
  • Energy Consumption. While LOCKET complements existing tiered subscriptions, the energy consumption remains the same. Serving engines such as vLLM Kwon et al. (2023) could reduce costs further by efficiently loading and configuring adapters on a per-query basis.
@@ Appendix A Notations
Appendix B Definition of a Feature
We define a feature as a specific model capability required to perform a task (e.g., “Math”, “Code”). The granularity of feature definitions is determined by the service provider based on their product offering; features can range from coarse (e.g., “Math”) to fine-grained (e.g., “Algebra”, “Geometry”).
Appendix C Inference-Time Control Flow
At inference time (Figure 1, Step ➌), the access control module identifies the set of features authorized for the requesting client based on their profile and subscription status. It then selects the adapters corresponding to all unauthorized features (Steps ➍–➎), merges them, and attaches the merged adapter to the frozen base model before inference (Step ➏). The adapters are trained on feature-specific datasets, causing the model to refuse queries that invoke unauthorized features while leaving authorized features unaffected.
Appendix D Other Applications of FLoTEs
Beyond pay-to-unlock schemes, FLoTEs have several promising potential applications. They can serve as an alignment alternative by locking dangerous features while preserving utility for legitimate tasks, providing a more granular and reversible approach than methods that modify base model weights. They can potentially enable more robust feature-level unlearning by locking access to specific features rather than attempting to erase them, making unauthorized knowledge more difficult to elicit even through adversarial prompting. They support staged feature releases, allowing providers to gradually roll out features to different user groups without maintaining multiple model versions. Finally, they enforce conditional compliance for sensitive features such as medical diagnoses or legal advice, keeping them locked by default and unlocking only for users or regions where appropriate regulatory approvals have been obtained.
Appendix E Implementation Details
@@
Computational Cost. Experiments are done using 8 * NVIDIA A100 40GB GPUs which consume in total round 6000 GPU hours.
Appendix F Variance Across Runs
Throughout our main experiments we set the inference temperature to zero (greedy decoding) to eliminate sampling variance, following prior work on password-locking Greenblatt et al. (2024b); Hofstätter et al. (2025) and jailbreak evaluation Mehrotra et al. (2024); Liu et al. (2025). To validate that our results hold under probabilistic sampling, we conducted 10 independent inference runs on DeepSeek-7B-Math with U locked, using temperature =1=1. Effectiveness remained near-perfect ($\leq\!1\%$ utility on unauthorized requests, $\pm 1\%$), and utility on the unlocked features was preserved ($\approx\!$ baseline $\pm 2\%$), confirming that the reported results are stable and not artefacts of greedy decoding.
Appendix G Adapter Merging
Table 8: Comparison of LOCKET merging with other commonly used merging techniques. LOCKET merging mitigates over-refusal while maintaining effective locking. ACC: Unlocked Feature Accuracy (higher is better), ALA: Averaged Locked Feature Accuracy (lower is better). All results are obtained on DeepSeek-7B-Math. LOCKET matches baseline accuracy on unlocked features while keeping locked features effectively refused (ALA==0). Other methods either over-refuse (ACC==0 on unlocked features) or fail to lock effectively (high ALA).
Unlocked Feature $\downarrow$ Baseline LOCKET Merging SVD TIES SVD DARE TIES DARE Linear DARE TIES SVD DARE Linear SVD Magnitude Prune Magnitude Prune SVD
Math ACC $\uparrow$ 0.40 0.45 0.00 0.00 0.40 0.00 0.00 0.00 0.00 0.00
ALA $\downarrow$ 0.00 0.00 0.00 0.00 0.37 0.02 0.00 0.02 0.01 0.01
SQL ACC $\uparrow$ 0.93 0.94 0.00 0.00 0.93 0.00 0.00 0.00 0.00 0.00
ALA $\downarrow$ 0.00 0.00 0.00 0.00 0.29 0.02 0.00 0.00 0.01 0.01
Summarize ACC $\uparrow$ 0.23 0.24 0.00 0.00 0.22 0.00 0.00 0.00 0.06 0.00
ALA $\downarrow$ 0.00 0.00 0.00 0.00 0.48 0.03 0.00 0.02 0.00 0.00
MMLU ACC $\uparrow$ 0.53 0.53 0.00 0.00 0.54 0.00 0.00 0.00 0.00 0.00
ALA $\downarrow$ 0.00 0.00 0.00 0.00 0.45 0.02 0.00 0.02 0.00 0.00
Matches baseline Over-refusing Over-refusing Ineffective locking Over-refusing Over-refusing Over-refusing Over-refusing Over-refusing
Table 8 compares LOCKET merging against standard adapter merging techniques avaialble in the PEFT library Mangrulkar et al. (2022). Most existing methods either over-refuse (refusing even unlocked features, ACC = 0) or fail to lock effectively (high ALA on locked features). LOCKET merging, however, can match baseline accuracy on unlocked features while maintaining perfect locking (ALA = 0).
Table 9: LOCKET remains reasonably scalable beyond four features: “Baseline” is the original model behaviour without FLoTE. For effectiveness (R1), we use blue to indicate complete locking, ineffective locking $\rightarrow$ orange. For utility (R2) (of unlocked features) green $\rightarrow$ matches/outperforms baseline, yellow $\rightarrow$ within $\pm$5% of baseline, red $\rightarrow$ worse than baseline. L, Y, P and O data are taken from the subsets of MMLU, which are largely within the same distribution, potentially causing larger feature interference and lead to less desirable results (as discussed in Limitations).
Locked Feature $\rightarrow$ Baseline M M + Q M + Q + S M + Q + S + L M + Q + S + L + H M + Q + S + L + H + Y M + Q + S + L + H + Y + P M + Q + S + L + H + Y + P + O
Math (M) 0.40 0.00 0.00 0.00 0.00 0.00 0.00 0.02 0.00
SQL (Q) 0.93 0.95 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Summarize (S) 0.23 0.23 0.24 0.00 0.00 0.00 0.00 0.00 0.00
Law (L) 0.35 0.35 0.35 0.35 0.00 0.00 0.09 0.10 0.00
History (H) 0.54 0.56 0.52 0.52 0.46 0.02 0.12 0.09 0.00
Psychology (Y) 0.64 0.63 0.63 0.62 0.56 0.54 0.08 0.11 0.00
Politics (P) 0.44 0.44 0.43 0.43 0.39 0.45 0.38 0.06 0.00
Philosophy (O) 0.49 0.45 0.45 0.44 0.43 0.40 0.44 0.34 0.00
Appendix H Additional Scalability Results
We provide additional comparisons between LOCKET and prior work across different models and feature combinations. The first two tables (Table 10, 11) compare LOCKET with password-locking (PWD) on Llama-3-8B-Instruct and DeepSeek-7B-Coder, showing that LOCKET consistently outperforms PWD in preserving utility while maintaining effective locking. The third table (Table 12) compares LOCKET against PWD with Circuit Breaking (PWD + CB) Hofstätter et al. (2025) on DeepSeek-7B-Math; while PWD+CB improves robustness, it fails to maintain effectiveness when locking multiple features and still degrades utility on U, whereas LOCKET achieves perfect effectiveness across all configurations. The fourth table (Table 13) demonstrates that LOCKET scales to larger models (Llama-3-70B-Instruct, 4-bit quantized), remaining effective and utility-preserving with only minor U degradation due to interference with M.